OpenStudy (pphalke):

Suppose the sales tax rate is 9% and a store is offering a 15% discount. What is the total cost (including sales tax) of a taxable item that has a list price of $700?

OpenStudy (layshay12):

So you first have to figure out the price of the product with the sales tax because they apply that first

OpenStudy (pphalke):

how do I figure that out

OpenStudy (layshay12):

You would have to do 700 times 9%. Which would give you?

OpenStudy (pphalke):

63

OpenStudy (layshay12):

Then you add that to your original price. now you take your new price of $763 and divid that by 15%

OpenStudy (layshay12):

Then you subtract that number from your 763 and you'll have your discounted price

OpenStudy (pphalke):

I get a 432 answer is it right? @layshay12

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

doesn't seem right to me

OpenStudy (misty1212):

try \[700\times 1.09\times .85\] with a calculator

OpenStudy (misty1212):

you do not get 432

OpenStudy (pphalke):

@misty1212 hey its 648.55

OpenStudy (layshay12):

That works too @misty1212 but i was never taught that way

OpenStudy (misty1212):

yes it is

OpenStudy (layshay12):

That's your answer

OpenStudy (misty1212):

btw the commutative property shows that it doesn't matter what you do first, because \[1.09\times .85=.85\times 1.09\]

OpenStudy (layshay12):

Exactly. It's just where I go to school we have a different style of learning curriculum